OpenStudy (palwandew):

A parabolic arch has a 100 meter base and height of 25 meter. Find the height of the arch at the point 30 meter from the centre of the base.

OpenStudy (janu16):

is this pre calc?

OpenStudy (palwandew):

Analytical Geometry

OpenStudy (welshfella):

write the parabola in vertex from the vertex is at (0,25) and the parabola opens downwards so we have y = -a(x - b)^2 + c where (b,c) is the vertex y = -a(x - 0)^2 + 25

OpenStudy (welshfella):

negative a because it opens downwards

OpenStudy (welshfella):

Find a by plugging in another point on the graph ( 50,0) so you can complete the equation of the parabola. Then find the required height by plugging in x = 30 and finding y.

OpenStudy (welshfella):

ok?

OpenStudy (palwandew):

Okay thanks.

OpenStudy (welshfella):

yw